# Influence of Exchange-Correlation Functionals and Neural Network Architectures on Li+-Ion Conductivity in Solid-State Electrolyte from Molecular Dynamics Simulations with Machine-Learning Force Fields

### v1

@@ Abstract
With the rapid advancement of machine learning techniques for materials’ simulations, machine-learned force fields (MLFFs) have emerged as a powerful tool that complements first-principles calculations by enabling high-accuracy molecular dynamics simulations over extended timescales. Typically, MLFFs are trained on data generated from density functional theory (DFT) using a specific exchange–correlation (XC) functional, with the goal of reproducing materials’ properties predicted by DFT with high fidelity. However, the uncertainties in MLFF-based simulations — arising from variations in both MLFF model architectures and the choice of XC functionals — are often not well understood.
@@
In this work, we construct MLFF models of different architectures parametrized on DFT data from both semilocal and hybrid functionals to describe Li+ diffusion in the solid-state electrolyte Li6PS5Cl. The influence of different exchange–correlation functionals on the Li+ diffusion coefficient is systematically investigated. To reduce statistical uncertainty, the mean squared displacements are averaged over 300 independent molecular dynamics (MD) trajectories of 70 ps each. The statistical variation of the extracted diffusion coefficients is below $1\%$, enabling an unambiguous assessment of the influence on the Li+ diffusivity from the exchange-correlation functional and from the MLFF models. Due to its tendency to underestimate band gaps and migration barriers, the semilocal functional predicts consistently higher Li+ diffusion coefficients, compared to the hybrid functional. Furthermore, comparisons among various neural network methods reveal that the differences in predicted diffusion coefficients arising from different network architectures are of the same order of magnitude as those caused by different functionals, indicating that the choice of the network model itself substantially influences the MLFF predictions. This observation calls from an urgent need for standardized protocols to minimize model-dependent biases in MLFF-based MD.
keywords:
machine-learned force fields , exchange-correlation functionals, neural network architectures , ion migration
1 Introduction
@@
With the continuous advancement of artificial intelligence (AI), MD simulations have entered a new stage. The rapid development of MLFFs has enabled MD simulations to retain the accuracy of first-principles calculations while achieving the computational efficiency of classical force fields [miwa2017interatomic, zhang2018deep, bartok2010gaussian, satorras2021n, xie2024gptff]. Increasing evidence shows that force field models based on neural networks exhibit excellent generalization ability, robust stability, and good transferability [noe2020machine, behler2007generalized, huang2021deep, li2025mechanistic, wu2021deep, gomez2024neural, li2025gas, shi2025soliton]. With the progress of deep learning theory and hardware resources, the amount of training data required has been significantly reduced, training efficiency has improved, and the model architectures have evolved from simple fully connected networks to more sophisticated and efficient designs such as graph neural networks and equivariant neural networks. Remarkably, by learning from higher-level electronic structure theories, MLFFs can even surpass the accuracy of conventional approximations of DFT. For instance, Liu et al. [liu2022phase] proposed a $\Delta$ -learning approach to construct force fields with random phase approximation (RPA) accuracy [ren2012random, chen2017random], while Di et al. [fan2025unexpected] employed the DeePMD-kit framework to develop force fields based on both PBE and HSE06 functionals to investigate the thermodynamic stability and phase transitions of HfO2.
@@
To address these issues, we examine three representative MLFF approaches—DeePMD-kit (DP) [zhang2018deep, zhang2018end], NequIP [batzner20223], and MACE [batatia2022mace]—combined with datasets generated using the popular XC functionals, the Perdew-Burke-Ernzerhof (PBE)[perdew1996generalized] generalized gradient approximation (GGA) and the Heyd-Scuseria-Ernzerhof (HSE) [heyd2003hybrid, paier2006screened, Ren/etal:2013, Gao/Guo:2011] screened hybrid density functionals (HDF). By comparing the predictive performance of different MLFF models in estimating the ionic conductivity of the solid-state electrolyte system—Li6PS5Cl(LPSC)[deiseroth2008li6ps5x, boulineau2012mechanochemical], we systematically investigate how the choices of MLFF architectures affect the prediction of physical properties and the biases they may introduce. Through this comparative analysis, we further evaluate the reliability and limitations of current MLFF methodologies for applications in battery materials.
2 Results
2.1 Consistency of Training Accuracy in MLFFs
@@
In this work, we start by training six machine-learning models to obtain accurate interatomic potentials for the LPSC system. The primary requirement for training a high-quality MLFF is to have a large and accurate dataset from first-principles calculations. Based on these accurate datasets, six MLFFs were trained using three neural network architectures: DP [zhang2018deep, zhang2018end], NequIP [batzner20223], and MACE [batatia2022mace]. The MACE framework provides a two-stage training scheme, in which the second training stage can significantly reduce the energy errors. Due to the computational cost associated with HSE calculations, the test-set errors were evaluated only for the PBE datasets. The root-mean-square errors (RMSEs) in energies and forces for the trained MLFFs are summarized in Table. 1. Detailed procedures for generating and selecting the training and test sets can be found in the 4. Methods section.
@@
The essential aspect of the present research is to ensure that the trained MLFFs are sufficiently accurate. Table. 1 demonstrates that the energy and force errors of all force-field models are rather small. Furthermore, the close agreement between the PBE MLFF predictions and the corresponding DFT reference values for the training and test set configurations (Fig. 1) confirms the accuracy of the constructed force fields.
@@
All three neural network models successfully reduce energy errors to the meV/atom scale. Among them, DP achieves the lowest energy RMSE, albeit with higher force errors. In contrast, the equivariant neural networks(ENN) NequIP and MACE achieve significantly lower force errors, while sacrificing some accuracy in energy, consistent with previous reports [batzner20223, batatia2022mace].
To investigate the influence of different neural network architectures and XC functionals on MLFFs in terms of the ionic conductivity of solid-state electrolytes, it is crucial to minimize the statistical errors in the estimation of diffusion coefficients. Here, the migration behavior of Li+ ions in LPSC was studied using MLFF-based MD simulations. The procedure is consistent with the AIMD approach: the mean squared displacement (MSD) of Li+ ions is computed from MD trajectories, and the diffusion coefficient $D$ is extracted via the Einstein relation [he2018statistical, mo2012first],
@@
Here, $d=3$ denotes the dimensionality of the system, and $N$ is the number of mobile ions. The angular brackets $\langle\cdots\rangle$ indicate averaging over multiple initial times $t$, which is essential for obtaining statistically meaningful MSD values.
@@
2.2 Statistical Error of Diffusion Coefficients
To extract diffusion coefficients faithfully from MD simulations, the statistical errors in evaluating the MSD (Eq. 1) need to be rigorously assessed and controlled. Obviously, the number of diffusion events captured within picosecond-scale simulations is often limited—particularly at low temperatures—leading to poor statistical accuracy in evaluating diffusion properties. Therefore, sufficiently long simulation times are required to suppress statistical errors [he2018statistical, lin2024comparative]. To obtain reliable diffusion coefficients across different temperatures, we employed a strategy of performing multiple short MD simulations, each initialized with random velocities but starting from the same structure [ong2015materials]. This approach maximizes sampling and mitigates statistical uncertainty in diffusion events. Specifically, in the temperature range of 700–1200 K, 300 independent trajectories of 70 ps each were generated using different MLFFs. For each trajectory, the MSD was first computed (Fig. 3a); then an average over all trajectories was performed, and the averaged MSD is indicated by the thick black line in Fig. 3a. Although the MSDs derived from individual trajectories show large variations in their slopes and can deviate significantly from linear behavior, the averaged MSD curve is much less prone to such statistical uncertainties.
@@ 2.2 Statistical Error of Diffusion Coefficients
2.3 Influence of the XC Functional on MLFF Performance
The choice of the XC functional significantly affects the accuracy of DFT calculations. While PBE offers computational efficiency, it notoriously underestimates band gaps, whereas hybrid functionals such as HSE provide improved electronic structure descriptions [banerjee2025theoretical]. For LPSC, PBE and HSE yield a band gap difference of 1.356 eV (Fig. 4). HDFs typically yield a higher barrier height than semilocal functionals [Ren/etal:2013, Gao/Guo:2011], which are expected to give rise to a smaller diffusion coefficient.
@@ 2.3 Influence of the XC Functional on MLFF Performance
2.4 Influence of Neural Network on MLFFs
@@ 3 Discussion
4 Methods
In this work, the Atomic-orbital Based Ab-initio Computation at UStc (ABACUS) [chen2010systematically, li2016large, lin2024abacus, zhou2025ABACUS] software package is used to create energies and forces encompassing diverse chemical environments in LPSC. Specifically, the PBE[perdew1996generalized] GGA and the HSE[heyd2003hybrid] screened HDF are adopted as the XC functionals in DFT calculations. A linear combination of numerical atomic orbital (NAO) basis sets at the double-zeta plus polarization (DZP) level, as developed in Ref. [Lin2021PRB], is employed in all calculations. Furthermore, the multi-projector “SG15-ONCV”-type norm-conserving pseudopotentials are employed to describe the interactions between nuclear ions and valence electrons. The efficient HSE implementation in ABACUS [lin2020accuracy, lin2020efficient, lin2025efficient] allows us to generate adequate training data for this system at the level of hybrid functionals.
In the present work, self-consistent calculations was converged within a threshold of $10^{-8}$ eV, and a $3\times 3\times 3$ k-mesh was adopted for the conventional unit cell. To model the atomic environments under stress, the lattice was stretched and compressed by up to $5\%$ relative to its equilibrium configuration. DFT-based data collection was carried out in two parts. The first part involved performing 10 ps of AIMD simulations on the unit cell (serving as the initial training set) in the NvT ensemble, with a time step of 1 fs. These simulations were conducted at temperatures ranging from 800 K to 1200 K, in increments of 200 K. The second part comprised static DFT calculations on structures selected during each DP-GEN iteration [zhang2019active, zhang2020dp]; these structures were generated via MLFF-driven MD simulations in both NvT and NpT ensembles, over a temperature range of 200 K to 1200 K. Through the iterative DP-GEN workflow, four DP force-field models of comparable accuracy were obtained (Fig. A1). Among them, the DP-00 model-showing the lowest energy and force RMSE-was selected for subsequent MD simulations.
Next, using the $\it{Simplify}$ module in DP-GEN, hundreds of thousands of atomic configurations sampled between 200–1200 K were screened, resulting in 283 representative structures. An additional 60 configurations were randomly selected from the remaining dataset as the test set. Static DFT calculations were then performed on these structures using both the PBE and HSE functionals, yielding two training datasets. Based on these datasets, six MLFFs were constructed using three neural network architectures: DP[zhang2018deep, zhang2018end], NequIP [batzner20223], and MACE [batatia2022mace]. The NequIP and MACE models were trained on the 283 selected configurations, while the DP-HSE model was obtained by transferring the DP-PBE model to the HSE-evaluated dataset.
Appendix A. Model-to-Model Variability in DeePMD-kit
@@
Acknowledgment
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0500201), and by the National Natural Science Foundation of China (Grants Nos. 52172258, 12134012, 12374067, and 12188101). This work was also funded by the National Key Research and Development Program of China (Grant Nos. 2022YFA1403800 and 2023YFA1507004) and by the robotic AI-Scientist platform of Chinese Academy of Sciences. The numerical calculations in this study were partly carried out on the ORISE Supercomputer, and partly on the SunRising-1 computing platform. This computing resource was also provided by the Bohrium Cloud Platform (https://bohrium.dp.tech), which is supported by DP Technology.

### v2

@@ Abstract
With the rapid advancement of machine learning techniques for materials’ simulations, machine-learned force fields (MLFFs) have emerged as a powerful tool that enables high-accuracy molecular dynamics (MD) simulations over extended timescales. Typically, MLFFs are trained on data generated from density functional theory (DFT) using a specific exchange–correlation (XC) functional, with the goal of reproducing materials’ properties predicted by DFT with high fidelity. However, the uncertainties in MLFF-based simulations — arising from variations in both MLFF model architectures and the choice of XC functionals — are often not well understood.
@@
In this work, we construct MLFF models of different architectures trained on DFT data with both semilocal and hybrid functionals to describe Li+ diffusion in the solid-state electrolyte Li6PS5Cl. The influence of different XC functionals on the Li+ diffusion coefficient is systematically investigated. To reduce statistical uncertainty, the mean squared displacements are averaged over 300 independent MD trajectories of 70 ps each. The statistical variation of the extracted diffusion coefficients is below $1\%$, enabling an unambiguous assessment of the influence on the Li+ diffusivity from the XC functional and the MLFF models. Due to its tendency to underestimate band gaps and migration barriers, the semilocal functional predicts higher Li+ diffusion coefficients, compared to the hybrid functional. Furthermore, comparisons across MLFF models reveal that differences arising from neural network architecture are as significant as, or even greater than, those from the choice of XC functional. This observation highlights an urgent need for standardized protocols to minimize model-dependent biases in MLFF-based MD.
I Introduction
@@
With the continuous advancement of artificial intelligence (AI), MD simulations have entered a new stage. The rapid development of MLFFs has enabled MD simulations to retain the accuracy of first-principles calculations while achieving the computational efficiency of classical force fields Miwa and Ohno (2017); Zhang et al. (2018a); Bartók et al. (2010); Satorras et al. (2021, 2022); Xie et al. (2024). Increasing evidence shows that force field models based on neural networks exhibit excellent generalization ability, robust stability, and good transferability Noé et al. (2020); Behler and Parrinello (2007); Huang et al. (2021); Li et al. (2025a); Wu et al. (2021); Gomez et al. (2024); Li et al. (2025b); Shi et al. (2025). With the progress of deep learning theory and hardware resources, the amount of training data required has been significantly reduced, training efficiency has improved, and the model architectures have evolved from simple fully connected networks to more sophisticated and efficient designs such as graph neural networks and equivariant neural networks. Remarkably, by learning from higher-level electronic structure theories, MLFFs can even surpass the accuracy of conventional approximations of DFT. For instance, Liu et al. Liu et al. (2022) proposed a $\Delta$ -learning approach to construct force fields with random phase approximation (RPA) accuracy Ren et al. (2012); Chen et al. (2017), while Di et al. Fan et al. (2025) employed the DeePMD-kit framework to develop force fields based on both PBE and HSE06 functionals to investigate the thermodynamic stability and phase transitions of HfO2.
@@
To address these issues, we examine three representative MLFF approaches—DeePMD-kit (DP) Zhang et al. (2018a, b), NequIP Batzner et al. (2022), and MACE Batatia et al. (2022)—combined with datasets generated using the popular XC functionals, the Perdew-Burke-Ernzerhof (PBE)Perdew et al. (1996) generalized gradient approximation (GGA) and the Heyd-Scuseria-Ernzerhof (HSE) Heyd et al. (2003); Paier et al. (2006); Ren et al. (2013); Gao and Guo (2011) screened hybrid density functionals (HDF). By comparing the predictive performance of different MLFF models in estimating the ionic conductivity of the solid-state electrolyte system—Li6PS5Cl(LPSC)Deiseroth et al. (2008); Boulineau et al. (2012), we systematically investigate how the choices of MLFF architectures affect the prediction of physical properties and the biases they may introduce. Through this comparative analysis, we further evaluate the reliability and limitations of current MLFF methodologies for applications in battery materials.
II Computational details
II.1 Preparation of First-principles Data and The training of MLFFs
We start by training six MLFF models to obtain accurate interatomic potentials for the LPSC system based on three neural-network architectures and data from two XC functionals. The primary requirement for training a high-quality MLFF is to have a large and accurate dataset from first-principles calculations. To this end, the Atomic-orbital Based Ab-initio Computation at UStc (ABACUS) Chen et al. (2010); Li et al. (2016); Lin et al. (2024b); Zhou et al. (2025) software package is used to create energies and forces encompassing diverse chemical environments in LPSC. Specifically, the PBEPerdew et al. (1996) GGA and the HSEHeyd et al. (2003) screened HDF are adopted as the XC functionals in DFT calculations. A linear combination of numerical atomic orbital (NAO) basis sets at the double-zeta plus polarization (DZP) level, as developed in Ref. Lin et al. (2021), is employed in all calculations. Furthermore, the multi-projector “SG15-ONCV”-type norm-conserving pseudopotentials are employed to describe the interactions between nuclear ions and valence electrons. The efficient HSE implementation in ABACUS Lin et al. (2020a, b, 2025) allows us to generate adequate training data for this system at the level of hybrid functionals.
In the present work, self-consistent calculations were converged within an accuracy of $10^{-8}$ eV, and a $3\times 3\times 3$ k-mesh was adopted for the conventional unit cell. To model the variation of atomic environments under stress, the lattice was stretched and compressed by up to $5\%$ relative to its equilibrium geometry. DFT-based data collection was conducted in two stages. First, ab initio molecular dynamics (AIMD) simulations were performed on the unit cell in the NVT ensemble for 10 ps with a time step of 1 fs. These simulations were carried out at temperatures ranging from 800 K to 1200 K in increments of 200 K. The second stage consisted of static DFT calculations on structures selected during each iteration of the DP-GEN workflow Zhang et al. (2019, 2020). These structures were generated through MLFF-driven molecular dynamics simulations in both NVT and NpT ensembles across a temperature range of 200 K to 1200 K.
Next, using the $\it{Simplify}$ module in DP-GEN, hundreds of thousands of atomic configurations sampled between 200–1200 K were screened, resulting in 283 representative structures. An additional 60 configurations were randomly selected from the remaining dataset as the test set. Static DFT calculations were then performed on these structures using both the PBE and HSE functionals, yielding two training datasets.
@@ II.1 Preparation of First-principles Data and The training of MLFFs
Based on these accurate datasets, six MLFFs were trained using three neural network architectures: DP Zhang et al. (2018a, b), NequIP Batzner et al. (2022), and MACE Batatia et al. (2022). The NequIP and MACE models were trained on the 283 selected configurations, while the DP-HSE model was obtained by transferring the DP-PBE model to the HSE-evaluated dataset. The MACE framework provides a two-stage training scheme, in which the second training stage can significantly reduce the energy errors. The root-mean-square errors (RMSEs) in energies and forces for the trained MLFFs are summarized in Table 1.
@@
The essential aspect of the present research is to ensure that the trained MLFFs are sufficiently accurate. Table 1 demonstrates that the energy and force errors of all force-field models are rather small. Furthermore, the close agreement between the PBE MLFF predictions and the corresponding DFT reference values for the training and test sets (Fig. 1) confirms the accuracy of the constructed force fields.
@@
All three neural network models successfully reduce energy errors to the meV/atom scale. Among them, DP achieves the lowest energy RMSE, albeit with higher force errors. In contrast, the equivariant neural networks(ENN) NequIP and MACE achieve significantly lower errors in force, while sacrificing some accuracy in energy, consistent with previous reports Batzner et al. (2022); Batatia et al. (2022). Furthermore, through the iterative DP-GEN workflow, four DP force-field models of comparable accuracy were obtained; a detailed discussion of these additional models is provided in Appendix A. Among them, the model exhibiting the smallest RMSE in energies and forces was selected for subsequent MD simulations (as shown in Fig. A1).
II.2 Ion Conductivity from Molecular Dynamics
To investigate the influence of different neural network architectures and XC functionals on MLFFs in terms of the ionic conductivity of solid-state electrolytes, it is crucial to minimize the statistical errors in the estimation of diffusion coefficients. Here, the migration behavior of Li+ ions in LPSC was studied using MLFF-based MD simulations. The procedure is consistent with the AIMD approach: the mean squared displacement (MSD) of Li+ ions is computed from MD trajectories, and the diffusion coefficient $D$ is extracted via the Einstein relation He et al. (2018); Mo et al. (2012),
@@ II.2 Ion Conductivity from Molecular Dynamics
Here, $d=3$ denotes the dimensionality of the system, and $N$ is the number of mobile ions.
@@
III Results
III.1 Statistical Error of Diffusion Coefficients
To extract diffusion coefficients faithfully from MD simulations, the statistical errors in evaluating the MSD (Eq. 1) need to be rigorously assessed and controlled. Obviously, the number of diffusion events captured within picosecond-scale simulations is often limited—particularly at low temperatures—leading to poor statistical accuracy in evaluating diffusion properties. Therefore, sufficiently long simulation times are required to suppress statistical errors He et al. (2018); Lin et al. (2024a). To obtain reliable diffusion coefficients across different temperatures, we employed a strategy of performing multiple short MD simulations, each initialized with random velocities but starting from the same structure Ong et al. (2015). This approach maximizes sampling and mitigates statistical uncertainty in diffusion events. Specifically, in the temperature range of 700–1200 K, 300 independent trajectories of 70 ps each were generated using different MLFFs. For each trajectory, the MSD was first computed (Fig. 3a); then an average over all trajectories was performed, and the averaged MSD is indicated by the thick black line in Fig. 3a. Although the MSDs derived from individual trajectories show large variations in their slopes and can deviate significantly from linear behavior, the averaged MSD curve is much less prone to such statistical uncertainties.
@@ III.1 Statistical Error of Diffusion Coefficients
III.2 Influence of the XC Functional on MLFF Performance
The choice of the XC functional significantly affects the accuracy of DFT calculations. While PBE offers computational efficiency, it notoriously underestimates band gaps, whereas hybrid functionals such as HSE provide improved electronic structure descriptions Banerjee and Morris (2025). For LPSC, PBE and HSE yield a band gap difference of 1.356 eV (Fig. 4). HDFs typically yield a higher barrier height than semilocal functionals Ren et al. (2013); Gao and Guo (2011), which are expected to give rise to a smaller diffusion coefficient.
@@ III.2 Influence of the XC Functional on MLFF Performance
III.3 Influence of Neural Network on MLFFs
@@ IV Conclusion
Acknowledgment
This work was supported by the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0500201), and by the National Natural Science Foundation of China (Grants Nos. 52172258, 12134012, 12374067, and 12188101). This work was also funded by the National Key Research and Development Program of China (Grant Nos. 2022YFA1403800 and 2023YFA1507004) and by the robotic AI-Scientist platform of Chinese Academy of Sciences. The numerical calculations in this study were partly carried out on the ORISE Supercomputer. This computing resource was also provided by the Bohrium Cloud Platform (https://bohrium.dp.tech), which is supported by DP Technology.
Competing Interests
The authors declare no competing interests.
Appendix A Model-to-Model Variability in DeePMD-kit